OpenStudy (anonymous):

PLEASE HELP! Will become fan. And give medal. 16. The area of a rectangular table is given by the trinomial x2 + 7x – 30. What are the possible dimensions of the rectangle? Use factoring. (1 point) a. (x+10) and (x-3) b. (x-10) and (x-3) c. (x+10) and (x+3) d. (x-10) and (x+3)

OpenStudy (anonymous):

@agent0smith can you help at all? :/

OpenStudy (anonymous):

\[\huge x^2 + 7x – 30= x^2 + 10x-3x – 30\] \[\huge = x(x + 10)-3(x +10) = (x+10) (x-3) \] Hence option a. (x+10) (x-3) are the required dimensions of the rectangle. @anon99

OpenStudy (anonymous):

@dpasingh could you help with two more?

OpenStudy (anonymous):

Sure go ahead..But type the question in ask question by mentioning me.. @anon99

OpenStudy (anonymous):

okay :)